# Verifying LLM Inference to Prevent Model Weight Exfiltration

### v1

@@ Abstract
As large AI models become increasingly valuable assets, the risk of model weight exfiltration from inference servers grows accordingly. An attacker controlling an inference server may exfiltrate model weights by hiding them within ordinary model outputs, a strategy known as steganography. This work investigates how to verify model responses to defend against such attacks and, more broadly, to detect anomalous or buggy behavior during inference. We formalize model exfiltration as a security game, propose a verification framework that can provably mitigate steganographic exfiltration, and specify its minimal trust assumptions. To enable verification, we characterize valid sources of non-determinism in large language model inference and introduce two practical estimators for them. We evaluate our detection framework on several open-weight models ranging from 3B to 30B parameters. On MOE-Qwen-30B, our detector reduces exfiltratable information to $0.5\%$ with false-positive rate of $10^{-2}$, corresponding to a $200\times$ slowdown for adversaries. Overall, this work further establishes a foundation for defending against model weight exfiltration and demonstrates that strong protection can be achieved with minimal additional cost to inference providers.  [^1] Our code is made available publicly.
@@ 1 Introduction
Malicious adversaries can exploit many potential channels, from infiltrating development environments and stealing backups to abusing side channels. While conventional security mechanisms can mitigate some of these risks, defending the inference pipeline poses a unique challenge. To support normal operation, inference providers must allow vast volumes of data—often hundreds of gigabytes per day—to flow in and out of their data centers [NLK+24]. This open communication channel creates an unavoidable tension between accessibility and security.
@@
In this work, we introduce two methods for estimating the likelihood that each generated token was honestly produced by the claimed model (GM-LS and IPT-LS). This enables us to measure deviations from expected model behavior and derive information-theoretic bounds on the amount of information that even a capable adversary could exfiltrate through normal inference traffic. We propose a detection scheme that can be integrated into larger model security systems.
@@
    We propose the Fixed-Seed Sampling Likelihood (FSSL) as a method for estimating likelihood of being sampled, and we introduce two estimates for this, Gumbel-Max Likelihood Score (GM-LS), and IPT Likelihood Score (IPT-LS), whic can be used for LLMs that sample with the Gumbel-Max trick and Inverse Probability Transforms, respectively.
@@
    Validates empirically on production systems We evaluate exfiltration rates on Llama-3.1-8B, Llama-3.2-3B, Qwen1.5-MoE-7B, Qwen3-30B-A3B. For Qwen3-30B-A3B we demonstrate a $0.5\%$ exfiltrable information rate with $10^{-2}\%$ FPR — a ${\geq}200\times$ slowdown in attack. For a 1 TB model with 10001000Gb of traffic per day, this extends exfiltration time from ${\sim}1$ days to >200{>}200 days.
@@ 2.1 Basic Setting (LLM inference)
Transformer inference consists of two phases: prefill processes the input prompt in a single parallel forward pass, while decode autoregressively generates output tokens one at a time. Though both phases have similar asymptotic complexity, prefill achieves significantly higher throughput in practice (typically 3–10×) because it is compute-bound and highly parallelizable, whereas decode is memory-bandwidth bound due to repeatedly loading the key–value cache [Erd25]. This makes prefill particularly advantageous for verification or evaluation tasks that do not require autoregressive generation.
@@ 2.2 Sampling From Large Language Models
Sampling from a Large Language model can be thought of in two steps: 1. generating the probability distribution, and then 2. sampling a token from it. For step 1, Given logits $z\in\mathbb{R}^{V}$ over a vocabulary of size $V$, the model induces a categorical distribution $\pi_{i}=\textrm{softmax}(z)_{i}=\exp(z_{i})/\sum_{j=1}^{V}\exp(z_{j})$.. To sample from a probability distribution, many modern LLMs sample tokens using either the Inverse Probability Transform (e.g., Ollama [Oll], HuggingFace [WDS+20]) or the Gumbel-Max Trick (e.g., VLLM [KLZ+23]). After applying temperature scaling and filtering (e.g., top-$k$, top-$p$, or min-$p$), the model converts logits into probabilities via the softmax function.
@@
1:Seed $s$, probability vector $\mathbf{p}\in[0,1]^{V}$ with $\sum_{t=1}^{V}p_{t}=1$
@@
1:Seed $s$, probability vector $\vec{p}\in[0,1]^{V}$ (over tokens $1..V$, $\sum_{i}p_{i}=1$)
@@ 2.3 Non-determinism in Machine Learning
Fixing the random seed (even with temperature T=0T{=}0) does not guarantee identical outputs. Non-determinism arises because floating-point arithmetic is non-associative ($(a{+}b){+}c\neq a{+}(b{+}c)$): different GPU kernels or shapes can accumulate tensors in different orders, yielding different values. Batch variance, where kernel computation strategies depend on the batch size, which depends on the concurrent load, is the most common source of numerical noise. Inference optimizations further shift computation orders—batch size, attention chunking/prefix caching, and cache reuse—and MoE models add routing variability that can depend on load and tiny numerical perturbations [HL25].
Identical open-weight models have shown cross-provider performance discrepancies, consistent with differences in templates, quantization/serving stacks, and the variability sources above [Wil25].
@@
To illustrate non-determinism in machine learning, we run a simulation of this using the OpenAI API. Using a single fixed prompt and identical seeds/sampling parameters, we generate 1,0001{,}000 completions and build a token-level prefix tree. Each node $v$ represents a prefix $y_{1:t}$ and the fork factor $\phi(v)$ is the number of possible next-token. On OpenAI 4o-mini, we observe that tokens have dramatically varying levels of non-determinism; the vast majority of nodes have $\phi(v)=1$, with divergences concentrated at a small number of positions (Fig. 2).
(a) Token-level prefix tree for 1,0001{,}000 seeded completions (first 100100 positions shown). Red nodes mark fork points (nodes with >1>1 children).
@@ 2.4 Steganography and Watermarking
There is a recent but long line of work in language model watermarking [Aar22, KGW+23a, CGZ24, FGJ+25, KTHL23, CG24, GM25, CHS25]. However, most works operate in a model that require the watermarker and detector to share only a small key with the goal of making the watermark as robust as possible. This means that even the highest throughput schemes only embed about $1$ bit per “block,” where a block is text that is being watermarked [^3] The length of a “block” depends on the scheme and security parameter, but often requires about 128-bits of empirical entropy. [CHS25]. In contrast, language model steganography [KJGR21] has a higher throughput rate, since it assumes text is unchanged. These schemes can almost reach capacity, embedding as many bits as the entropy of the response.
In this work, we adopt the language of steganography, with minor modifications listed below. For other cryptographic objects, such as pseudorandom functions, hash functions, etc. we refer the reader to definitions provided in Katz and Lindell [KL14].
@@
Note that there are many other variants of steganography, including stateless, public-key, robust, etc. [AP02, HLVA02, VAH04]. However, in our context, we wish to give a potential exfiltrator as much theoretical power as possible. So we assume that both endpoints are stateful and can have an arbitrarily long secret key in common.
@@ 2.4.1 Relevant Results in Steganography
This theorem is an immediate consequence of Shannon’s source coding theorem [Sha48, S+59]. This theorem states: (i) that no random variable $X$ can be encoded with fewer bits than $H(X)$ in expectation and (ii) that there exists an encoding of $X$ with expected length at most $\lceil H(X)\rceil$. Therefore, if $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ and $\mathsf{Steg}$ is correct with probability $1$, then one could encode $\mathbf{M}$ as the covertext, using a random key $K$ and random state $\mathsf{st}$, $\mathbf{C}\leftarrow\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$ and decoded with $\mathsf{Dec}(K,\mathsf{st},\mathbf{C})$. Then, based on the source coding theorem, $\mathbf{C}\sim\mathcal{D}_{\mathsf{Steg}}$ (and therefore $\mathbf{M}$) can be encoded with $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ bits, which is a direct contradiction. ∎
In contrast to the above theorem and to illustrate the difficulty of preventing covert exfiltration, we recall a well-known theorem from [HLVA02] stated in Theorem 2.3. This shows that assuming classical cryptography exists at all, i.e., one-way functions exist, then one can construct a steganographic scheme that is near-optimal.
Theorem 2.3 (Due to [HLVA02]).
@@ 2.5 Model Weight Exfiltration
In the context of discussing model-weight exfiltration, a recent framework [NLK+24] introduces the notions of security levels from 1 to 5, which increase in defensive capacity. SL1 is a system that can likely thwart amateur attempts. SL2 is a system that can likely thwart most professional opportunistic efforts by attackers that execute moderate-effort or nontargeted attacks SL3 is a system that can likely thwart cybercrime syndicates or insider threats. SL4 is system that can likely thwart most standard operations by leading cyber-capable institutions. SL5 is a system that could plausibly be claimed to thwart most top-priority operations by the top cyber-capable institutions.
@@ 4.2 Solution 1: Perplexity and Syntactic Filters
Perplexity filtering. A slightly more sophisticated approach uses a reference language model trained on benign message traffic to compute sequence-level perplexity—a standard measure of how well a model predicts a sequence. Formally, perplexity is the exponential of the average negative log-likelihood assigned to each token; lower values indicate that the model finds the text highly predictable, while higher values suggest the text is surprising or out-of-distribution. Messages with unusually high or low perplexity—indicating atypical token statistics—are flagged as potential exfiltration attempts [Kar24]. This can detect outputs that are syntactically normal but statistically implausible under the benign distribution, though adaptive adversaries can still craft messages that maintain natural perplexity while encoding information.
@@ 4.4 Valid Non-determinism with a Fixed Seed: The Fixed-Seed Posterior Distribution
While Inverse-Probability Transform and Gumbel-Max Trick are mathematically deterministic, in practice the logit distribution $\vec{p}_{\theta}(x)$ is produced by a non-deterministic computation. Even with a fixed seed $\sigma$, floating-point imprecision, asynchronous GPU execution, and nondeterministic kernel scheduling can cause small but measurable differences in $\vec{p}_{\theta}(x)$ across runs. This effect arises even in benign settings, independent of any adversarial manipulation, and has been well-documented in prior work [YLD+25]. We refer to this inherent variability as valid non-determinism.
@@ 4.5 Solution 3: Classification using the Fixed-Seed Sampling Likelihood Policy
with $\tau$ calibrated on benign traffic to meet a target false–positive rate. Per-token decisions (or summed log-likelihoods) can be aggregated over sequences, users, or time windows for higher-level alerts. We instantiate estimates of the FSSL for for two samplers: Gumbel-Max Likelihood Score (GM-LS) for samplers that use the Gumbel-Max Trick and IPT Likelihood Score (IPT-LS) for samplers that use Inverse Probability Transforms for sampling; subsequent sections derive their estimators and evaluate performance.
@@ 4.6 Theoretical Exfiltration Limitations
    the seed is expanded by a cryptographic hash function, [^6] Formally, our bounds need the stronger assumption that we can model the hash function as a random oracle [BR93].
@@
Here we provide some estimates on how tight we expect the above bounds to be against a motivated adversary. Theorem 4.1 is believed to be tight in theory. As long as basic cryptography exists, steganography exists as a consequence of Theorem 2.3 [HLVA02]. Even in practice, an adversary that carefully designs a steganographic scheme against the FSSL policy should likely be able to get close to this bound. They could likely draw heavily on techniques from steganographic literature to design schemes.
@@ 5.1 The Cost of Verification versus Generation
During verification, a single token mismatch between two generations makes their subsequent probability distributions divergent, as they are conditioned on different contexts (e.g., all continuations of “the cat sat on” vs. “the cat sat beside” will diverge). Instead of continuing togenerating text with fixed seeds—which would diverge immediately—we exploit the fact that LLMs expose logits for all next-token predictions during the prefill stage. A single prefill over the full input–output sequence yields conditional probabilities for every position, enabling token-level suspiciousness scores in a single forward pass. Verification therefore involves only prefill, whereas normal inference requires both prefill and autoregressive decode. In short-context regimes, decoding achieves substantially lower efficiency than $3–5\times$ lower hardware utilization is typical for short-context decoding than prefill, which implies a $\penalty 10000\ 3–5\times$ throughput gap. This gap increases in long context regimes, although the exact difference will depend on attention optimizations and implementation details [Erd25].
5.2 Scoring Function: IPT Likelihood Score (IPT-LS)
In Section 4.5 we introduce the notion of a classification using Fixed-Seed Sampling Likelihood. In this section we introduce a technique for approximating the fixed-seed sampling likelihood (FSSL) for the inverse probability transform sampling technique (IPT). We introduce a method IPT Likelihood Score (IPT-LS). The core assumption is that across repeated model evaluations, the per-token probability vectors $\vec{p}_{\theta}(x)$ may vary due to small numerical or stochastic perturbations, but such variation is relatively small.

### v2

# Verifying LLM Inference to Detect Model Weight Exfiltration

@@ Abstract
As large AI models become increasingly valuable assets, the risk of model weight exfiltration from inference servers grows accordingly. An attacker controlling an inference server may exfiltrate model weights by hiding them within ordinary model responses, a strategy known as steganography. This work investigates how to verify LLM model inference to defend against such attacks and, more broadly, to detect anomalous or buggy behavior during inference. We formalize model weight exfiltration as a security game, propose a verification framework that can provably mitigate steganographic exfiltration, and specify the trust assumptions associated with our scheme. To enable verification, we characterize valid sources of non-determinism in large language model inference and introduce two practical estimators for them. We evaluate our detection framework on several open-weight models ranging from 3B to 30B parameters. On MOE-Qwen-30B, our detector reduces exfiltratable information to $0.5\%$ with false-positive rate of $10^{-2}$, corresponding to a $200\times$ slowdown for adversaries. Overall, this work further establishes a foundation for defending against model weight exfiltration and demonstrates that strong protection can be achieved with minimal additional cost to inference providers. Our code is made available at: github.com/RoyRin/inference_verification_for_model_weight_exfiltration.
@@ 1 Introduction
Malicious adversaries can exploit many potential channels, from infiltrating development environments and stealing backups to abusing side channels. While conventional security mechanisms can mitigate some of these risks, defending the inference pipeline poses a unique challenge. To support normal operation, inference providers must allow vast volumes of data—often hundreds of gigabytes per day—to flow in and out of their data centers [nevo2024securing]. This open communication channel creates an unavoidable tension between accessibility and security.
@@
In this work, we introduce two methods for estimating the likelihood that each generated token was honestly produced by the claimed model (Token-DiFRand Token-IPT-DiFR). This enables us to measure deviations from expected model behavior and derive information-theoretic bounds on the amount of information that even a capable adversary could exfiltrate through normal inference traffic. We propose a detection scheme that can be integrated into larger model security systems.
@@
    We propose the Fixed-Seed Sampling Likelihood (FSSL) as a method for estimating likelihood of being sampled, and we introduce two estimates for this, Token-DiFR (FSSL-GM), and Token-IPT-DiFR (FSSL-IPT), whic can be used for LLMs that sample with the Gumbel-Max trick and Inverse Probability Transforms, respectively.
@@
    Validates empirically on production systems We evaluate exfiltration rates on Llama-3.1-8B, Llama-3.2-3B, Qwen1.5-MoE-7B, Qwen3-30B-A3B. For Qwen3-30B-A3B we demonstrate a $0.5\%$ exfiltrable information rate with $10^{-2}\%$ FPR, a ${\geq}200\times$ slowdown in attack. For a 1 TB model with 10001000Gb of traffic per day, this extends exfiltration time from ${\sim}1$ days to >200{>}200 days.
@@ 2.1 Basic Setting (LLM inference)
Transformer inference consists of two phases: prefill processes the input prompt in a single parallel forward pass, while decode autoregressively generates output tokens one at a time. Though both phases have similar asymptotic complexity, prefill achieves significantly higher throughput in practice (typically 3–10×) because it is compute-bound and highly parallelizable, whereas decode is memory-bandwidth bound due to repeatedly loading the key–value cache [erdil2025inferenceeconomicslanguagemodels]. This makes prefill particularly advantageous for verification or evaluation tasks that do not require autoregressive generation.
@@ 2.2 Sampling From Large Language Models
Sampling from a Large Language model can be thought of in two steps: 1. generating the probability distribution, and then 2. sampling a token from it. For step 1, Given logits $z\in\mathbb{R}^{V}$ over a vocabulary of size $V$, the model induces a categorical distribution $\pi_{i}=\textrm{softmax}(z)_{i}=\exp(z_{i})/\sum_{j=1}^{V}\exp(z_{j})$.. To sample from a probability distribution, many modern LLMs sample tokens using either the Inverse Probability Transform (e.g., Ollama [ollama_docs], HuggingFace [huggingface]) or the Gumbel-Max Trick (e.g., VLLM [vllm]). After applying temperature scaling and filtering (e.g., top-$k$, top-$p$, or min-$p$), the model converts logits into probabilities via the softmax function.
@@
1:Seed $\sigma$, probability vector $\mathbf{p}\in[0,1]^{V}$ with $\sum_{t=1}^{V}p_{t}=1$
@@
1:Seed $\sigma$, probability vector $\vec{p}\in[0,1]^{V}$ (over tokens $1..V$, $\sum_{i}p_{i}=1$)
@@ 2.3 Non-determinism in Machine Learning
Fixing the random seed (even with temperature T=0T{=}0) does not guarantee identical outputs. Non-determinism arises because floating-point arithmetic is non-associative ($(a{+}b){+}c\neq a{+}(b{+}c)$): different GPU kernels or shapes can accumulate tensors in different orders, yielding different values. Batch variance, where kernel computation strategies depend on the batch size, which depends on the concurrent load, is the most common source of numerical noise. Inference optimizations further shift computation orders—batch size, attention chunking/prefix caching, and cache reuse—and MoE models add routing variability that can depend on load and tiny numerical perturbations [he2025nondeterminism].
Identical open-weight models have shown cross-provider performance discrepancies, consistent with differences in templates, quantization/serving stacks, and the variability sources above [willison_nondeterminism].
@@
To illustrate non-determinism in machine learning inference, we query the OpenAI API with a single fixed prompt, using identical seed and sampling parameters (temperature $0.7$, seed $1234$) across 20,00020{,}000 completions. We tokenize each completion (using whitespace) and construct a token-level prefix tree: each node $v$ represents a prefix $y_{1:t}$, and the fork factor $\phi(v)$ is the number of distinct next-tokens observed. On gpt-4o-mini, we observe that non-determinism is sparse but present: the vast majority of nodes have $\phi(v)=1$ (deterministic), while divergences concentrate at a small number of token positions (Fig. 2).
(a) Token-level prefix tree for 1,0001{,}000 seeded completions (first $100$ positions shown). Red nodes mark fork points (nodes with >1>1 children).
@@ 2.4 Steganography and Watermarking
There is a recent but long line of work in language model watermarking [Aar22, ICML:KGWKMG23, COLT:ChrGunZam24, CIC:FGJMMW23, ARXIV:KTHL23, pseudo_random_error_correcting, golowich2025editdistancerobustwatermarks, SP:CHS25]. However, most works operate in a model that require the watermarker and detector to share only a small key with the goal of making the watermark as robust as possible. This means that even the highest throughput schemes only embed about $1$ bit per “block,” where a block is text that is being watermarked [^2] The length of a “block” depends on the scheme and security parameter, but often requires about 128-bits of empirical entropy. [SP:CHS25]. In contrast, language model steganography [CCS:KJGR21] has a higher throughput rate, since it assumes text is unchanged. These schemes can almost reach capacity, embedding as many bits as the entropy of the response.
In this work, we adopt the language of steganography, with minor modifications listed below. For other cryptographic objects, such as pseudorandom functions, hash functions, etc. we refer the reader to definitions provided in Katz and Lindell [KatLin14].
@@
Note that there are many other variants of steganography, including stateless, public-key, robust, etc. [JOC:AP02, C:HLvA02, IDK:vAH04]. However, in our context, we wish to give a potential exfiltrator as much theoretical power as possible. So we assume that both endpoints are stateful and can have an arbitrarily long secret key in common.
@@ 2.4.1 Relevant Results in Steganography
This theorem is an immediate consequence of Shannon’s source coding theorem [shannon1948mathematical, shannon1959coding]. This theorem states: (i) that no random variable $X$ can be encoded with fewer bits than $H(X)$ in expectation and (ii) that there exists an encoding of $X$ with expected length at most $\lceil H(X)\rceil$. Therefore, if $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ and $\mathsf{Steg}$ is correct with probability $1$, then one could encode $\mathbf{M}$ as the covertext, using a random key $K$ and random state $\mathsf{st}$, $\mathbf{C}\leftarrow\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$ and decoded with $\mathsf{Dec}(K,\mathsf{st},\mathbf{C})$. Then, based on the source coding theorem, $\mathbf{C}\sim\mathcal{D}_{\mathsf{Steg}}$ (and therefore $\mathbf{M}$) can be encoded with $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ bits, which is a direct contradiction. ∎
In contrast to the above theorem and to illustrate the difficulty of preventing covert exfiltration, we recall a well-known theorem from [C:HLvA02] stated in Theorem 2.3. This shows that assuming classical cryptography exists at all, i.e., one-way functions exist, then one can construct a steganographic scheme that is near-optimal.
Theorem 2.3 (Due to [C:HLvA02]).
@@ 2.5 Model Weight Exfiltration
In the context of discussing model-weight exfiltration, a recent framework [nevo2024securing] introduces the notions of security levels from 1 to 5, which increase in defensive capacity. SL1 is a system that can likely thwart amateur attempts. SL2 is a system that can likely thwart most professional opportunistic efforts by attackers that execute moderate-effort or nontargeted attacks SL3 is a system that can likely thwart cybercrime syndicates or insider threats. SL4 is system that can likely thwart most standard operations by leading cyber-capable institutions. SL5 is a system that could plausibly be claimed to thwart most top-priority operations by the top cyber-capable institutions.
@@ 4.2 Solution 1: Perplexity and Syntactic Filters
Perplexity filtering. A slightly more sophisticated approach uses a reference language model trained on benign message traffic to compute sequence-level perplexity—a standard measure of how well a model predicts a sequence. Formally, perplexity is the exponential of the average negative log-likelihood assigned to each token; lower values indicate that the model finds the text highly predictable, while higher values suggest the text is surprising or out-of-distribution. Messages with unusually high or low perplexity—indicating atypical token statistics—are flagged as potential exfiltration attempts [perplexityfilter]. This can detect outputs that are syntactically normal but statistically implausible under the benign distribution, though adaptive adversaries can still craft messages that maintain natural perplexity while encoding information.
@@ 4.4 Valid Non-determinism with a Fixed Seed: The Fixed-Seed Posterior Distribution
While Inverse-Probability Transform and Gumbel-Max Trick are mathematically deterministic, in practice the logit distribution $\vec{p}_{\theta}(x)$ is produced by a non-deterministic computation. Even with a fixed seed $\sigma$, floating-point imprecision, asynchronous GPU execution, and nondeterministic kernel scheduling can cause small but measurable differences in $\vec{p}_{\theta}(x)$ across runs. This effect arises even in benign settings, independent of any adversarial manipulation, and has been well-documented in prior work [yuan2025fp32deathchallengessolutions]. We refer to this inherent variability as valid non-determinism.
@@ 4.5 Solution 3: Classification using the Fixed-Seed Sampling Likelihood Policy
with $\tau$ calibrated on benign traffic to meet a target false–positive rate. Per-token decisions (or summed log-likelihoods) can be aggregated over sequences, users, or time windows for higher-level alerts. We instantiate estimates of the FSSLfor two samplers: we introduce Token-DiFR to estimate FSSLquantity for samplers that use the Gumbel-Max Trick, FSSL-GM, and we introduce Token-IPT-DiFR (FSSL-IPT) for samplers that use Inverse Probability Transforms for sampling, which estiamtes FSSL-IPT. Subsequent sections derive their estimators and evaluate performance.
@@ 4.6 Theoretical Exfiltration Limitations
    the seed is expanded by a cryptographic hash function, [^5] Formally, our bounds need the stronger assumption that we can model the hash function as a random oracle [CCS:BelRog93].
@@
Here we provide some estimates on how tight we expect the above bounds to be against a motivated adversary. Theorem 4.1 is believed to be tight in theory. As long as basic cryptography exists, steganography exists as a consequence of Theorem 2.3 [C:HLvA02]. Even in practice, an adversary that carefully designs a steganographic scheme against the FSSL policy should likely be able to get close to this bound. They could likely draw heavily on techniques from steganographic literature to design schemes.
@@ 5.1 The Cost of Verification versus Generation
During verification, a single token mismatch between two generations makes their subsequent probability distributions divergent, as they are conditioned on different contexts (e.g., all continuations of “the cat sat on” vs. “the cat sat beside” will diverge). Instead of continuing togenerating text with fixed seeds—which would diverge immediately—we exploit the fact that LLMs expose logits for all next-token predictions during the prefill stage. A single prefill over the full input–output sequence yields conditional probabilities for every position, enabling token-level suspiciousness scores in a single forward pass. Verification therefore involves only prefill, whereas normal inference requires both prefill and autoregressive decode. In short-context regimes, decoding achieves substantially lower efficiency than $3–5\times$ lower hardware utilization is typical for short-context decoding than prefill, which implies a $\penalty 10000\ 3–5\times$ throughput gap. This gap increases in long context regimes, although the exact difference will depend on attention optimizations and implementation details [erdil2025inferenceeconomicslanguagemodels].
5.2 Inverse Probability Transform Scoring Function Token-IPT-DiFR
In Section 4.5 we introduce the notion of a classification using Fixed-Seed Sampling Likelihood. In this section we introduce a technique for approximating the fixed-seed sampling likelihood (FSSL) for the inverse probability transform sampling technique (IPT). We investigate a method Token-IPT-DiFR, which is concurrently introduced in parallel work [tokendifr]. The core assumption is that across repeated model evaluations, the per-token probability vectors $\vec{p}_{\theta}(x)$ may vary due to small numerical or stochastic perturbations, but such variation is relatively small.